OpenStudy (anonymous):

Find exact value of cos(7pi)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cos%287pi%29 use the website

OpenStudy (zehanz):

Remember cosx starts and ends with value 1: cos(0) = 1, cos(2pi)=1. Its period is 2pi. Max values at beginning and end, so min value in the middle: cos(pi) = -1. Because the period is 2pi, adding 2pi doesn't make a difference: you end up at the same height. So: cos(pi) = cos(3pi) = cos(5pi) = cos(7pi) = -1.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

approx =.9272

OpenStudy (anonymous):

-1